A boy is dropped from a height of 250 m, and simultaneously another body is dropped from a height of 200 m. What is the difference in heights after they have fallen for 5 seconds?

Grade:9

1 Answers

John Sharon Sandeep
6 years ago
If thought logically, they both accelerate at the same rate(9.8) and start with same initial velocity(0) and dropped at the same time. Hence difference in height after any number of seconds= 250m-200m=50m. If you need the answer the physics way, here it is-- u1=u2=0,t=5s,g=9.8ms-1➡by S=ut+1/2at2 , S1=0+9.8×5×5/2=122.5m & S2=0+9.8×5×5/2=122.5m. Therefore distance of boy from ground=250-122.5=127.5m & distance of object from ground=200-122.5=77.5m.🔨 Therefore difference between them=127.5-77.5=50m.